# Investigating VTubing as a Reconstruction of Streamer Self-Presentation: Identity, Performance, and Gender

### v1

@@ 5.3. Revisiting VTubing through Lens of Objectification Theory
The objectification theory posits that women are frequently sexually objectified via the male gaze (Mulvey, 1989) and that the prevalence of such objectification can lead to internalization, also known as self-objectification, where women internalise the male gaze as the primary view of themselves (Fredrickson and Roberts, 1997; Moradi and Huang, 2008). Self-objectification causes negative consequences such as body shaming, depression, anxiety, eating disorders, etc. (Moradi and Huang, 2008; Calogero, 2004).
@@
Unlike other social media or avatar-mediated systems, VTubing affords greater agency in gender-related customization and performance, where VTubers can infuse gender and sexuality into their avatars via outfit design, animation design, voicing, chatting with viewers, and so on. Such affordances open avenues for future investigation into the landscapes of gender expression by disentangling virtual and real gender identities. Future research can investigate motivations behind constructing specific gender identities, and the interplay between real and virtual gender. Particularly, we are interested in how female streamers consider their virtual gender customization before and after being a VTuber in a misogynist environment, and whether the “female dominance” of VTubing should be seen as an empowerment or objectification for women.
@@
Our research focused primarily on the customization and performance of the virtual gender alone. VTubing allowed for experimentation of new gender identities (I4) and gender expressions different from stereotypes (e.g., C5’s avatar design). Our interviewees also mentioned it alleviated the appearance anxiety of being a streamer and sometimes granted a sense of power (I3) and control over gender expression (I4). Nevertheless, such agency might not necessarily translate into empowered sexuality (Baumgardner and Richards, 2004; Attwood, 2007; Levy, 2010), especially when exposed to misogynistic viewers (Lu et al., 2021). Self-sexualization is prevalent and often used as a strategy to engage viewers (mostly male). It remains to be seen how such self-sexualization impacts psychological well-being of Nakanohitos, especially women, and whether it leads to self-objectification while VTubing via an often sexualized avatar that does not necessarily resemble themselves. We advocate for future efforts in CSCW to understand such issues through the lens of feminism. In particular, a survey study (e.g., using Twenty Statement Test (Fredrickson et al., 1998)) is needed to measure potential self-objectification in this scenario. Many interviewees explained that their motivations were self-driven and such sexualization was based solely on their virtual avatar. However, this is still concerning because previous studies in lab environments suggested that controlling a sexualized avatar that resembled one’s self might still lead to self-objectification in the real world (Fox et al., 2013, 2015).
@@
Despite sexualized gender expression, our interviewees almost unanimously reported less sexual harassment or hostility from viewers, which is uncommon in live streaming (Ruberg et al., 2019; Wotanis and McMillan, 2014; Döring and Mohseni, 2019; Freeman et al., 2015). The question of whether this phenomenon can be seen as empowerment also remains open for feminist debate. VTubing could potentially act as an expedient for women or members of other minority groups to stream, however, such practices are still limited, since they make no effort to change misogynistic culture (Lu et al., 2021; Ruberg et al., 2019), but in some ways, still caters to it for viewership. Similar to the concern of ‘raunch culture’ (Levy, 2010), the seeming female-dominance of VTubing that is entrenched in such a culture might actually imply a sense of false empowerment, as it works to maintain the sexual status quo.
@@ 5.4.2. Design for the Pseudonymous Identity
VTubers were found to value both the pseudonymity an avatar offers and the authenticity of their identity and performance. As prior work highlighted the sensitivity of Nakanohitos’ personal information, while calling for transparency of working conditions and fair treatment (Lu et al., 2021), we argue that the blurring line between the virtual and the real should be made more transparent and explicit to ensure the authenticity of VTubing and prevent VTubers from harassment. Further guidelines should be proposed in the VTuber community or by VTubing companies to inform new VTubers and viewers of related issues. For example, VTubers could specify ‘dos and don’ts’ for incoming viewers in their social media groups. VTubing companies can also promote ‘industry consensus’ to fan groups or new VTubers. Moreover, as inspired by (Wu et al., 2022), we call for VTubing platforms or software to better support privacy and identity management. We anticipate features that, for instance, remind viewers during live streams what content (either the performance or identity) might be virtual or authentic, as well as what should be considered private, as configured by VTubers themselves. Such practices are expected to reduce hostility or harassment and ensure an overall enjoyable and inclusive environment.

### v2

@@ 5.3. Revisiting VTubing through Lens of Objectification Theory
The objectification theory posits that women are frequently sexually objectified via the male gaze (Mulvey, 1989) and that the prevalence of such objectification can lead to internalization, also known as self-objectification, where women internalise the male gaze as the primary view of themselves (Fredrickson and Roberts, 1997; Moradi and Huang, 2008). Self-objectification causes negative consequences such as body shaming, depression, anxiety, eating disorders, etc. (Moradi and Huang, 2008; Calogero, 2004). Unlike other social media or avatar-mediated systems that often put women in objectified roles (e.g., (Davis, 2018)), VTubing does afford greater agency in gender-related design and performance, where VTubers, mostly women, can infuse customized gender and sexuality into their avatars via outfit design, animation design, voicing, chatting with viewers, and so on. Such affordances open avenues for future investigation into the landscapes of gender expression by disentangling virtual and real gender identities. Particularly, we are interested in how female streamers consider their virtual gender customization before and after being a VTuber in a misogynist environment, and whether the “female dominance” of VTubing should be seen as an empowerment or objectification for women.
@@
Our research focused primarily on the customization and performance of the virtual gender alone. VTubing allowed for experimentation of new gender identities (I4) and gender expressions different from stereotypes (e.g., C5’s avatar design). Our interviewees also mentioned it alleviated the appearance anxiety of being a streamer and sometimes granted a sense of power (I3) and control over gender expression (I4). Nevertheless, such agency might not necessarily translate into empowered sexuality (Baumgardner and Richards, 2004; Attwood, 2007; Levy, 2010), especially when exposed to misogynistic viewers (Lu et al., 2021b). Self-sexualization is prevalent and often used as a strategy to engage viewers (mostly male). It remains to be seen how such self-sexualization impacts psychological well-being of Nakanohitos, especially women, and whether it leads to self-objectification while VTubing via an often sexualized avatar that does not necessarily resemble themselves. We advocate for future efforts in CSCW to understand such issues through the lens of feminism. In particular, a survey study (e.g., using Twenty Statement Test (Fredrickson et al., 1998)) is needed to measure potential self-objectification in this scenario. Many interviewees explained that their motivations were self-driven and such sexualization was based solely on their virtual avatar. However, this is still concerning because previous studies in lab environments suggested that controlling a sexualized avatar that resembled one’s self might still lead to self-objectification in the real world (Fox et al., 2013, 2015).
@@
Despite sexualized gender expression, our interviewees almost unanimously reported less sexual harassment or hostility from viewers, which is uncommon in live streaming (Ruberg et al., 2019; Wotanis and McMillan, 2014; Döring and Mohseni, 2019). The question of whether this phenomenon can be seen as empowerment also remains open for feminist debate. VTubing could potentially act as an expedient for women or members of other minority groups to stream, however, such practices are still limited, since they make no effort to change misogynistic culture (Lu et al., 2021b; Ruberg et al., 2019), but in some ways, still caters to it for viewership. Similar to the concern of ‘raunch culture’ (Levy, 2010), the seeming female-dominance of VTubing that is entrenched in such a culture might actually imply a sense of false empowerment, as it works to maintain the sexual status quo.
@@ 5.4.2. Design for the Pseudonymous Identity
VTubers were found to value both the pseudonymity an avatar offers and the authenticity of their identity and performance. As prior work highlighted the sensitivity of Nakanohitos’ personal information, while calling for transparency of working conditions and fair treatment (Lu et al., 2021b), we argue that the blurring line between the virtual and the real should be made more transparent and explicit to ensure the authenticity of VTubing and prevent VTubers from harassment. Further guidelines should be proposed in the VTuber community or by VTubing companies to inform new VTubers and viewers of related issues. For example, VTubers could specify ‘dos and don’ts’ for incoming viewers in their social media groups. VTubing companies can also promote ‘industry consensus’ to fan groups or new VTubers. Moreover, as inspired by (Wu et al., 2022), we call for VTubing platforms or software to better support privacy and identity management. We anticipate features that, for instance, remind viewers during live streams what content (either the performance or identity) might be virtual or authentic, as well as what should be considered private, as configured by VTubers themselves. Such practices are expected to reduce hostility or harassment and ensure an overall enjoyable and inclusive environment.
@@ 6. Conclusion
Acknowledgements
This research was partially supported by the 2021 CCF-Tencent Rhino-Bird Research Fund and the Research Matching Grant Scheme (RMGS, Project No. 9229095). We thank Tencent, China Computer Federation (CCF), and Research Grants Council of Hong Kong for their support.